# Can weak-gravity, causality-violation arguments constrain modified gravity?

### v1

@@ V.1 Shapiro time delay for GR black holes
To see how the result of the last section intersects with the standard Shapiro time delay, we start by revisiting the computation for a non-rotating, Schwarzchild black hole in GR, following mostly [63]. The line element of a Schwarzschild black hole can be written as
@@
[…]
@@
[…]
$=\Delta\rho\left[1+\frac{2GM}{\Delta\rho}\ln\left.\left(\frac{\rho+\sqrt{\rho^{2}+b^{2}}}{b}\right)\right|_{\rho_{1}}^{\rho_{2}}\right.$
$\left.+{\cal{O}}\left(\frac{G^{2}M^{2}}{b^{2}}\right)\right].$ (41)
The second term on the second line of the above equation is a delay relative to the travel time of the null geodesic as measured by asymptotic observers, $\Delta\rho$. This is the standard Shapiro time delay. We emphasize that the weak-field expansion carried out in Eq. (40) is only for convenience; the exact result is in Eq. (39), which can be trivially expanded to any order in $GM/b$ that one desires.
In order to make contact with the results of the previous section, were obtained with scattering amplitude methods, we will consider the difference between time delays for two different impact parameters $b_{0}$ and $b$, with $b_{0}>b$. The largest delay is associated with the smallest impact parameter, so we define
$\Delta t=\delta_{b}t-\delta_{b_{0}}t\approx 2GM\ln\left(\frac{b_{0}}{b}\right)\,,$ (42)
@@
$\delta_{b}t$ $\simeq(\delta_{b}t)_{\rm Schw}-\left.\frac{a}{b}\frac{2GM}{\sqrt{\rho^{2}+b^{2}}}\rho\right|_{\rho_{1}}^{\rho_{2}}+{\cal{O}}\left(\frac{a^{2}}{b^{2}},\frac{G^{2}M^{2}}{b^{2}}\right),$
$=\Delta\rho\left[1+\frac{2GM}{\Delta\rho}\ln\left.\left(\frac{\rho+\sqrt{\rho^{2}+b^{2}}}{b}\right)\right|_{\rho_{1}}^{\rho_{2}}\right.$
@@
where $(\delta_{b}t)_{\rm Schw}$ is the travel time of a null ray that passes a Schwarzschild black hole in Eq. (V.1).
Let us discuss the above result in a bit more detail. The first thing we notice is that the spin correction to the travel time is a time advance, and not a time delay, i.e. the term proportional to $a$ in Eq. (51) is strictly negative. Of course, this time advance contribution does not violate causality because the Kerr spin parameter cannot take arbitrarily large values. One could demand that the term proportional to $a$ be smaller than the second term in Eq. (51) to find a “causality constraint” on the spin of black holes. This, of course, would yield a constraint that (i) would be outside the regime of validity of the Kerr solution (i.e. when $a>M$, the Kerr black hole contains a naked singularity), (ii) would be outside the regime of the slow-rotation approximation used to derive this constraint. The term proportional to $a$ in Eq. (51) was derived assuming that it would lead to a small correction to the Schwarzschild black hole case (because of the small spin expansion in Eq. (49)), and thus, no “causality constraint” on spin is actually possible.
@@ V.2 Shapiro time delay for dCS gravity black holes
Then, the leading-order-in-$\zeta$ contribution to the time delay in GR of Eq. (51), $\delta_{b}t\to\delta_{b}t+\delta_{b}^{\rm dCS}t$, is
[…]
@@
$+{\cal{O}}\left(\frac{G^{5}M^{5}}{b^{5}}\right)\,,$
$\sim\frac{a\zeta}{8}\left(\frac{GM}{b}\right)^{4}\left.\left(\frac{b}{\rho}\right)^{5}\right|_{\rho_{1}}^{\rho_{2}}+{\cal{O}}\left(\frac{G^{5}M^{5}}{b^{5}}\right)\,.$ (55)
Observe that the dCS correction is extremely suppressed by both $\zeta$ and $(GM/b)^{4}$. Since both $\zeta\ll 1$ (to remain within the regime of validy of dCS gravity as an effective field theory) and $GM/b\ll 1$ (to remain within the far-field expansion of the time delay calculation), this correction can never dominate over the GR contribution in Eq. (V.1).
Is this result consistent with the scattering amplitude calculation? Yes. The regime $\zeta<1$ corresponds to the limit $\ell_{\rm CS}<r_{s}$, for which the amplitude approach also does not yield a time advance. We could have carried out the above calculation without expanding in $GM/b\ll 1$, for example by considering light rays that graze the black hole horizon. Had we done so, however, the suppression of this term with respect to $\zeta$ would still be present, even after including nonlinearities that can take factors $GM/b\sim 1$ into account.
@@ V.3 Shapiro time delay from stars in dCS gravity
$\delta_{b}^{\rm dCS}t$ $=\int_{\rho_{1}}^{\rho_{2}}\left\{-2\frac{GM}{r}\frac{ab}{r^{2}}-\frac{128\pi}{3}\frac{ab}{(GM)^{2}}\frac{\zeta}{2}\left[C_{1}\left(\frac{GM}{r}\right)^{9}\right.\right.$
@@ VI.1 Time delays from shockwaves in GR
The energy-momentum tensor for a massless source propagating in the $-x$ direction (that is, moving along $u=0$) is
@@
to linear order in $h_{ij}$. Recall that the ${i,j}$ indices only run through directions transverse to $x$. Focusing on the $ij$ equation, we have
$02\partial_{u}\partial_{v}h_{ij}+H(u,x^{i})\partial^{2}_{v}h_{ij}-\partial_{k}\partial^{k}h_{ij}=$ 0 . (71)
@@ VI.2 Time delays from shockwaves in dCS gravity
$\delta C^{\mu\nu}$ =¯$=\overline{\nabla}_{\gamma}\phi\bar{\epsilon}^{\gamma\delta\alpha(\mu}\left(\frac{h}{2}\overline{\nabla}_{\alpha}\overline{R}^{\nu)}_{\;\;\delta}+\overline{\nabla}_{\alpha}\delta R^{\nu)}_{\;\;\delta}+\delta\Gamma^{\nu)}_{\alpha\lambda}\overline{R}^{\lambda}_{\;\;\delta}-\overline{R}^{\nu)}_{\;\;\lambda}\delta\Gamma^{\lambda}_{\alpha\delta}\right)-\frac{1}{2}\nabla_{\gamma}\phi\delta\Gamma^{\gamma}_{\lambda\delta}\epsilon^{\alpha\beta\lambda(\mu}\overline{R}^{\nu)\delta}_{\;\;\;\;\alpha\beta}+$
@@
To find the discontinuity in $u$, we evaluate this equation on a trajectory along the $u$-direction that passes the shock with an impact parameter $b$ in the transverse directions. Note that $b>|p|^{-1}$ for the consistency of the optical (eikonal) approximation of the wavepacket description to be valid. In that limit, $\partial_{u}^{2}\phi$ is subleading compared to $\partial_{u}\phi$, and we find
$-2\partial_{u}\partial_{v}h_{ij}-H\partial^{2}_{v}h_{ij}+4\tilde{\alpha}\kappa^{2}\partial_{u}\phi\epsilon^{uvk}_{\;\;\;\;\;\;(i}\partial_{|u}\partial^{2}_{v|}h_{j)k}=0.$ (88)
@@
where we defined $[h]^{+}_{-}=h(u=0^{+})-h(u=0^{-})$. Decomposing the metric perturbation in circular polarizations, $h_{ij}=e^{A}_{ij}h_{A}$, and using $\epsilon^{uvk}_{\;\;\;\;\;\;i}e^{ij}_{B}=\epsilon_{B}e_{B}^{kj}$ (no sum over $B$), gives us
$[h_{B}]^{+}_{-}=\left(-\frac{f(b)}{2}+4\tilde{\alpha}\kappa^{2}\sqrt{\frac{|p|}{b^{D-1}}}\epsilon_{B}\right)\left.\partial_{v}h_{B}\right|_{u=0},$ (92)
@@ VII Discussion and Conclusion
Consider dCS gravity as an example [see Eq. (IV.2)]. It comes with an energy scale $1/\sqrt{\alpha}$, and since it is a perturbatively nonrenormalizable gravitational theory, we expect higher-derivative terms to be suppressed by higher powers of $\alpha/\kappa$. However, due to the new $\phi hh$ cubic vertex, new diagrams start to significantly contribute to gravitational scattering at energy scales $\Lambda_{\rm dCS}\sim(M_{\rm Pl}/\alpha)^{1/3}$, where $\alpha$ is the coupling constant of the theory [68]. For $1/\sqrt{\alpha}<M_{\rm Pl}$, we have $\Lambda_{\rm dCS}>1/\sqrt{\alpha}>\kappa/\alpha$. Therefore, $\kappa/\alpha$ is the cutoff for processes away from the eikonal regime, especially for subplanckian scattering [^5] We thank Gregory Gabadadze for discussions about the dCS cutoff scale.. Moreover, as discussed in Sec. IV, to neglect graviton nonlinearities and remain within the validity of the eikonal approximation, we need $b\gg(\alpha r_{\rm s})^{1/3}$ or [see Eq. (33)],
@@ Acknowledgements.
SA thanks Dan Kabat for discussions. HB thanks Simon Caron-Huot for discussions on eikonal scattering. We also thank Robert Brandenberger, Pablo A. Cano, Sera Cremonini, Cyril Creque-Sarbinowski, Gregory Gabadadze, Francesco Serra, Javi Serra, Enrico Trincherini, and Leonardo G. Trombetta for comments on a earlier version of this work. We acknowledge support from the Simons Foundation through Award No. 896696. NY also acknowledges support from the NSF through award PHY-2207650 and NASA through Grant No. 80NSSC22K0806.

### v2

@@ V.1 Shapiro time delay for GR black holes
To see how the result of the last section intersects with the standard Shapiro time delay, we start by revisiting the computation for a non-rotating, Schwarzschild black hole in GR, following mostly [63]. The line element of a Schwarzschild black hole can be written as
@@
$\delta_{b}t$ $=\int_{\rho_{1}}^{\rho_{2}}\frac{\left(1-\frac{2GM}{r(\rho)}\frac{b^{2}}{r^{2}(\rho)}\right)^{1/2}}{\left(1-\frac{2GM}{r(\rho)}\right)}d\rho\,.$ (39)
@@
$\delta_{b}t$ $\simeq\int_{\rho_{1}}^{\rho_{2}}\left[1+\frac{2GM}{\sqrt{\rho^{2}+b^{2}}}+{\cal{O}}\left(\frac{G^{2}M^{2}}{b^{2}}\right)\right]d\rho$
$=\Delta\rho\left[1+\frac{GM}{\Delta\rho}\ln\left(\frac{\sqrt{\rho^{2}+b^{2}}+\rho}{\sqrt{\rho^{2}+b^{2}}-\rho}\right)\right|_{\rho_{1}}^{\rho_{2}}$
$\left.+{\cal{O}}\left(\frac{G^{2}M^{2}}{b^{2}}\right)\right]$
$=\Delta\rho\left.\left[1+\frac{2GM}{\Delta\rho}\ln\left(\frac{2\rho}{b}\right)\right|_{\rho_{1}}^{\rho_{2}}+{\cal{O}}\left(\frac{G^{2}M^{2}}{b^{2}},\frac{b^{2}}{\rho^{2}}\right)\right],$ (41)
where we assumed that $\rho_{1,2}\gg b$ in the last equality. The second term on the second line of the above equation is a delay relative to the travel time of the null geodesic as measured by asymptotic observers, $\Delta\rho$. This is the standard Shapiro time delay, which we can also write in terms of the radial coordinate of the emitter and receiver events $r_{e}\equiv r(\rho_{1})$ and $r_{o}\equiv r(\rho_{2})$, respectively (e.g., see Eq. (10.101) in [64]) We emphasize that the weak-field expansion carried out in Eq. (40) is only for convenience; the exact result is in Eq. (39), which can be trivially expanded to any order in $GM/b$ that one desires.
In order to make contact with the results of the previous section, which were obtained with scattering amplitude methods, we will consider the difference between time delays for two different impact parameters $b_{0}$ and $b$, with $b_{0}>b$. The largest delay is associated with the smallest impact parameter, so we define
$\Delta t\equiv\delta_{b}t-\delta_{b_{0}}t\approx 2GM\ln\left(\frac{b_{0}}{b}\right)\,$, (42)
@@
$(\delta_{b}t)_{\rm Kerr}$ $\simeq(\delta_{b}t)_{\rm Schw}-\left.\frac{a}{b}\frac{2GM}{\sqrt{\rho^{2}+b^{2}}}\rho\right|_{\rho_{1}}^{\rho_{2}}+{\cal{O}}\left(\frac{a^{2}}{b^{2}},\frac{G^{2}M^{2}}{b^{2}}\right),$
$=\Delta\rho\left[1+\frac{2GM}{\Delta\rho}\ln\left.\left(\frac{\sqrt{\rho^{2}+b^{2}}+\rho}{\sqrt{\rho^{2}+b^{2}}-\rho}\right)\right|_{\rho_{1}}^{\rho_{2}}\right$.
@@
where $(\delta_{b}t)_{\rm Schw}$ is the travel time of a null ray that passes a Schwarzschild black hole in Eq. (V.1). The analog of (42) after assuming that $\rho\gg b$ is
$(\Delta t)_{\rm Kerr}$ $=(\delta_{b}t)_{\rm Kerr}-(\delta_{b_{0}}t)_{\rm Kerr}$
$\approx 2GM\ln\left(\frac{b_{0}}{b}\right)-2GM\;\frac{a}{b}\left(1-\frac{b}{b_{0}}\right).$ (52)
Reproducing such a result from a scattering amplitude calculation would require having an incoming state with non-vanishing angular momentum.
Let us discuss the above results in a bit more detail. The first thing we notice is that the spin correction to the travel time is a time advance, and not a time delay, i.e. the term proportional to $a$ in Eqs. (51) and (52) is strictly negative because $b/b_{0}<1$ by convention. Of course, this time advance contribution does not violate causality because the Kerr spin parameter cannot take arbitrarily large values. One could demand that the term proportional to $a$ be smaller than the second term in Eq. (51) or in Eq. (52) to find a “causality constraint” on the spin of black holes. This, of course, would yield a constraint that (i) would be outside the regime of validity of the Kerr solution (i.e. when $a>M$, the Kerr black hole contains a naked singularity), (ii) would be outside the regime of the slow-rotation approximation used to derive this constraint. The term proportional to $a$ in Eqs. (51) and (52) was derived assuming that it would lead to a small correction to the Schwarzschild black hole case (because of the small spin expansion in Eq. (49)), and thus, no “causality constraint” on spin is actually possible.
@@ V.2 Shapiro time delay for dCS gravity black holes
Then, the leading-order-in-$\zeta$ contribution to the time delay in GR of Eq. (51), $(\delta_{b}t)_{\rm Kerr}\to(\delta_{b}t)_{\rm Kerr}+\delta_{b}^{\rm dCS}t$, is
$\delta_{b}^{\rm dCS}t$ $=\int_{\rho_{1}}^{\rho_{2}}\left[-\frac{5}{8}\zeta\left(\frac{GM}{r(\rho)}\right)^{4}\frac{ab}{r^{2}(\rho)}+{\cal{O}}\left(\frac{G^{5}M^{5}}{r^{5}}\right)\right]d\rho$
@@
$+{\cal{O}}\left(\frac{G^{5}M^{5}}{b^{5}}\right)$
$\approx\frac{a\zeta}{8}\left(\frac{GM}{b}\right)^{4}\left.\left(\frac{b}{\rho}\right)^{5}\right|_{\rho_{1}}^{\rho_{2}}\,.$ (56)
Observe that the dCS correction is extremely suppressed by both $\zeta$ and $(GM/b)^{4}$. Since both $\zeta\ll 1$ (to remain within the regime of validity of dCS gravity as an effective field theory) and $GM/b\ll 1$ (to remain within the far-field expansion of the time delay calculation), this correction can never dominate over the GR contribution in Eq. (V.1).
Is this result consistent with the scattering amplitude calculation? Yes. The regime $\zeta<1$ corresponds to the limit $\ell_{\rm CS}<r_{s}$, for which the amplitude approach also does not yield a time advance. We could have carried out the above calculation without expanding in $GM/b\ll 1$, for example by considering light rays that graze the black hole horizon. Had we done so, however, the suppression of this term with respect to $\zeta$ would still be present, even after including nonlinearities that can take factors of $GM/b\sim 1$ into account.
The dCS contribution to the total difference between time delays for impact parameters $b$ and $b_{0}>b$ is
$(\Delta t)_{\rm dCS}$ $\equiv\delta_{b}^{\rm dCS}t-\delta_{b_{0}}^{\rm dCS}t\approx-\frac{\zeta}{8}GM\left(\frac{b_{0}}{GM}\right)\left(\frac{a}{GM}\right)$
$\times\left(\frac{GM}{\rho_{1}}\right)^{5}\left(1-\frac{b}{b_{0}}\right)\left[1-\left(\frac{\rho_{1}}{\rho_{2}}\right)^{5}\right].$ (57)
As $\rho_{1}$ and $\rho_{2}$ originated from the integral along the null worldline parametrized by $\rho$, we have that $\rho_{2}>\rho_{1}$. Hence, the dCS correction to the time delay is always negative. This, however, does not mean that Eq. (57) leads to a time advance instead of a time delay, because this correction must be added on top of the Kerr time delay of Eq. (52). Only when $(\Delta t)_{\rm dCS}$ in Eq. (57) becomes larger than $(\Delta t)_{\rm Kerr}$ in Eq. (52) does one obtain a time advance. Note, however, that the dCS correction is suppressed relative to the Kerr correction by multiple terms: $\zeta\ll 1$ (so that the theory is within regime of validity), $a/(GM)\ll 1$ (so that the black hole solution be valid in the small spin approximation), $(GM/\rho_{1})^{5}\ll 1$ (so that the weak-field expansion of gravity be valid). Only when these assumptions, which the above calculations rely on, are broken can the dCS correction lead to a time advance.
The above results cannot and should not be compared directly with Eq. (28). This is because the calculation in this section computes a different quantity (one that is actually observable) than that computed with the scattering amplitude calculation. In this subsection, we were concerned with the time delay of a photon that travels through the spacetime of a black hole in dCS gravity. In the scattering amplitude calculation, one is concerned with the time delay between a wave in the scalar field that is sourced by gravitational perturbations relative to a wave in the gravitational perturbations. Since the scalar field in dCS gravity does not couple directly to matter, there is no way to directly observe the scalar wave, and thus, to measure this time difference. Nonetheless, regardless of the situation considered, we see that the only way dCS gravity can lead to a time advance is if one considers the theory outside of its regime of validity (ie. large $\zeta$) and outside the weak gravity regime.
@@ V.3 Shapiro time delay from stars in dCS gravity
[…]
@@ VI.1 Time delays from shock waves in GR
The energy-momentum tensor for a massless source propagating in the x-x direction (that is, moving along $u=0$) is
@@
to linear order in $h_{ij}$. Recall that the i,j{i,j} indices only run through directions transverse to $x$. Focusing on the $ij$ equation, we have
$2\partial_{u}\partial_{v}h_{ij}+H(u,x^{i})\partial^{2}_{v}h_{ij}-\partial_{k}\partial^{k}h_{ij}=0.$ (73)
@@ VI.2 Time delays from shock waves in dCS gravity
$\delta C^{\mu\nu}$ $=\overline{\nabla}_{\gamma}\phi\bar{\epsilon}^{\gamma\delta\alpha(\mu}\left(\frac{h}{2}\overline{\nabla}_{\alpha}\overline{R}^{\nu)}_{\;\;\delta}+\overline{\nabla}_{\alpha}\delta R^{\nu)}_{\;\;\delta}+\delta\Gamma^{\nu)}_{\alpha\lambda}\overline{R}^{\lambda}_{\;\;\delta}-\overline{R}^{\nu)}_{\;\;\lambda}\delta\Gamma^{\lambda}_{\alpha\delta}\right)-\frac{1}{2}\nabla_{\gamma}\phi\delta\Gamma^{\gamma}_{\lambda\delta}\epsilon^{\alpha\beta\lambda(\mu}\overline{R}^{\nu)\delta}_{\;\;\;\;\alpha\beta}+$
@@
To find the discontinuity in $u$, we evaluate this equation on a trajectory along the $u$-direction that passes the shock with an impact parameter $b$ in the transverse directions. Note that $b>|p|^{-1}$ for the consistency of the optical (eikonal) approximation of the wave-packet description to be valid. In that limit, $\partial_{u}^{2}\phi$ is subleading compared to $\partial_{u}\phi$, and we find
$2\partial_{u}\partial_{v}h_{ij}-H\partial^{2}_{v}h_{ij}+4\tilde{\alpha}\kappa^{2}\partial_{u}\phi\epsilon^{uvk}_{\;\;\;\;\;\;(i}\partial_{|u}\partial^{2}_{v|}h_{j)k}=0.$ (90)
@@
where we defined $h]^{+}_{-}=h(u=0^{+})-h(u=0^{-})$. Decomposing the metric perturbation in circular polarizations, $h_{ij}=e^{A}_{ij}h_{A}$, and using $\epsilon^{uvk}_{\;\;\;\;\;\;i}e^{ij}_{B}=\epsilon_{B}e_{B}^{kj}$ (no sum over $B$), gives us
$h_{B}]^{+}_{-}=\left(-\frac{f(b)}{2}+4\tilde{\alpha}\kappa^{2}\sqrt{\frac{|p|}{b^{D-1}}}\epsilon_{B}\right)\left.\partial_{v}h_{B}\right|_{u=0},$ (94)
@@ VII Discussion and Conclusion
Consider dCS gravity as an example [see Eq. (IV.2)]. It comes with an energy scale $1/\sqrt{\alpha}$, and since it is a perturbatively non-renormalizable gravitational theory, we expect higher-derivative terms to be suppressed by higher powers of $\alpha/\kappa$. However, due to the new $\phi hh$ cubic vertex, new diagrams start to significantly contribute to gravitational scattering at energy scales $\Lambda_{\rm dCS}\sim(M_{\rm Pl}/\alpha)^{1/3}$, where $\alpha$ is the coupling constant of the theory [69]. For $1/\sqrt{\alpha}<M_{\rm Pl}$, we have $\Lambda_{\rm dCS}>1/\sqrt{\alpha}>\kappa/\alpha$. Therefore, $\kappa/\alpha$ is the cutoff for processes away from the eikonal regime, especially for subplanckian scattering [^6] We thank Gregory Gabadadze for discussions about the dCS cutoff scale.. Moreover, as discussed in Sec. IV, to neglect graviton nonlinearities and remain within the validity of the eikonal approximation, we need $b\gg(\alpha r_{\rm s})^{1/3}$ or [see Eq. (33)],
@@ Acknowledgements.
SA thanks Dan Kabat for discussions. HB thanks Simon Caron-Huot for discussions on eikonal scattering. We also thank Robert Brandenberger, Pablo A. Cano, Simon Caron-Huot, Sera Cremonini, Cyril Creque-Sarbinowski, Gregory Gabadadze, Francesco Serra, Javi Serra, Enrico Trincherini, and Leonardo G. Trombetta for comments on a earlier version of this work. We acknowledge support from the Simons Foundation through Award No. 896696. NY also acknowledges support from the NSF through award PHY-2207650 and NASA through Grant No. 80NSSC22K0806.